OpenStudy (anonymous):

8(z-3)=2(z+12) Is this the right answer? z=3 and z=12

OpenStudy (anonymous):

you can get only one answer

OpenStudy (anonymous):

Ok what did something wrong?

OpenStudy (anonymous):

You should start by distributing the brackets. \[a(b-c)=a*b+a*(-c)\]

OpenStudy (anonymous):

What would be the correct way to solve this?

OpenStudy (anonymous):

?

OpenStudy (amistre64):

expand the terms and collect them

OpenStudy (anonymous):

If \[a(b-c)=a*b+a*(-c)\]then\[8(z-3)=?\]and\[2(z+12)=?\]

OpenStudy (anonymous):

So will I multiply 8*z and 8*3

OpenStudy (amistre64):

yep, and same thing to the other side

OpenStudy (anonymous):

Bingo! But remember that the 3 is \(negative\), so the product of 8 and -3 will also be negative.

OpenStudy (anonymous):

so What do I do to the answers on both sides?

OpenStudy (anonymous):

So am I adding 8z and 2z together?